# Approximate Bayesian Computation with Statistical Distances for Model Selection

### v3

@@ Abstract
62F15,
62-08,
62C10,
approximate Bayesian computation,
model choice,
animal behavior,
@@ 1 Introduction
While deep learning has been applied to related problems such as kinetic model classification (Burés and Larrosa, 2023), such tasks differ from true model selection, especially in overlapping model spaces. ABC, by simulating under each model and comparing to observed data, offers a principled framework for likelihood-free model comparison. In this paper, we evaluate the performance of full data ABC methods for model selection, comparing their effectiveness to summary-based ABC and deep learning classifiers. Our results demonstrate that full data ABC offers a flexible and robust approach to model choice when likelihoods are unavailable.
@@ 4 Simulation Study
The neural network trains on $10^{6}$ datasets (80% for training, 20% for testing) generated using the true parameters, assuming the true model is included in the simulation design. In contrast, ABC simulates from prior distributions and selects samples closest to the observed data using distance metrics, making it more flexible and less assumption-dependent, especially when the true model or parameters are unknown.
@@ 4.1 Normal Mean Hypothesis Test
Synthetic datasets are generated from the true model $\mathcal{N}(\theta_{0},1)$ with $\theta_{0}\in\{0.0,0.1,0.2,0.3,0.4,0.5\}$, allowing evaluation of method performance as the true mean departs from $\tilde{\theta}=0$. For each value of $\theta_{0}$, 100 datasets are simulated and analyzed using the seven methods described above for model selection and parameter inference. Under $H_{1}$, the prior is $\theta\sim N(\tilde{\theta},100)$, providing a weakly informative prior; under $H_{0}$, $\theta$ is fixed at $\tilde{\theta}=0$.
@@
We first assess each method’s ability to correctly identify the data-generating model. Performance is summarized using the estimated probability of selecting $H_{0}$ across different values of $\theta_{0}$ (Table 1), and confusion matrices and boxplots of posterior probabilities (Appendix D of the Supplementary Material). ABC methods based on distance metrics correctly identify $H_{0}$ in all cases, indicating good control of false positives. The transition is gradual: for small deviations from zero (e.g., $\theta_{0}=0.1$), the classifier shows limited power, but confidence in selecting $H_{1}$ grows as the true mean increases. When $\theta_{0}\geq 0.4$, $H_{1}$ is selected most of the time, reaching around 95% selection accuracy at $\theta_{0}=0.5$. ABC-SA follows a similar pattern but tends to select $H_{0}$ more often than the full data ABC approaches when the true mean increases. ABC-QDA selects $H_{0}$ in nearly all cases, while NN shows approximately random model selection, with probabilities close to 50% for both hypotheses.
We further evaluate posterior inference quality for $\theta$ by computing the MSE between the posterior mean and the true value $\theta_{0}$ across repetitions (Table 1). For each dataset, the posterior mean is computed from the model selected as most probable by each method. Results show that MSE is generally low, with the smallest values observed when $\theta_{0}=0$ and $\theta_{0}=0.5$, consistent with improved classification accuracy for these values. When the true mean is small but non-zero, posterior distributions for all methods retain substantial mass near zero, reflecting limited information in the data to distinguish between models. ABC-SA and ABC-QDA exhibit larger MSE than the distance-based ABC methods, while NN estimation is further affected by the difficulty in correctly identifying the model.
@@
These results demonstrate that ABC with distance metrics provides reliable performance for both model selection and parameter estimation, comparable to ABC-Stat using sufficient summary statistics. The comparison with other ABC algorithms and with NN in this simple example highlights the limitations of those methods and illustrates the advantages of full data ABC approaches for model selection problems.
@@
$\theta_{0}=0.0$ $\Pr(\theta=0)$ 0.975 0.976 0.975 0.973 0.500 0.976 0.981
$\hat{\theta}$ 0.000 0.000 0.000 0.002 0.049 0.000 0.000
@@
$\theta_{0}=0.1$ $\Pr(\theta=0)$ 0.931 0.933 0.929 0.921 0.503 0.948 0.974
$\hat{\theta}$ 0.007 0.006 0.007 0.010 0.049 0.000 0.000
@@
$\theta_{0}=0.2$ $\Pr(\theta=0)$ 0.788 0.796 0.789 0.760 0.495 0.847 0.955
$\hat{\theta}$ 0.046 0.042 0.042 0.048 0.104 0.010 0.000
MSE 0.037 0.037 0.038 0.038 0.009 0.037 0.040
$\theta_{0}=0.3$ $\Pr(\theta=0)$ 0.522 0.543 0.519 0.459 0.499 0.671 0.934
$\hat{\theta}$ 0.147 0.144 0.149 0.168 0.149 0.141 0.000
MSE 0.062 0.061 0.060 0.056 0.023 0.096 0.090
$\theta_{0}=0.4$ $\Pr(\theta=0)$ 0.210 0.236 0.203 0.157 0.503 0.450 0.902
$\hat{\theta}$ 0.344 0.335 0.345 0.364 0.000 0.285 0.000
MSE 0.039 0.041 0.038 0.034 0.160 0.125 0.160
$\theta_{0}=0.5$ $\Pr(\theta=0)$ 0.056 0.067 0.050 0.031 0.501 0.263 0.870
$\hat{\theta}$ 0.494 0.489 0.497 0.506 0.000 0.359 0.000
MSE 0.022 0.025 0.020 0.016 0.250 0.125 0.250
@@ 4.2 Exponential Family Model Selection
NN is the method that show the lowest accuracy of model selection: it correctly identifies the exponential model in only 78% of cases, the lognormal in 81%, and the gamma in around 20%. The gamma model was often misclassified as exponential, suggesting a structural limitation in distinguishing these distributions.
Parameter estimation results (Table 2) mirror these findings. For distance-based ABC methods, MSEs are consistently low when the correct model is selected, comparable to ABC-Stat. By contrast, NN produces biased estimates, reflecting its difficulty in identifying the correct model. ABC-SA yields slightly higher estimation errors. Finally, ABC-QDA has the largest MSE values across all experiments and ABC methods.
@@
$\mathcal{E}\text{xp}(0.5)$ $\Pr(M=M_{1})$ 0.883 0.836 0.953 0.850 0.948 0.952 0.775 0.772 0.922
$\hat{\theta}$ 0.512 0.506 0.511 0.515 0.514 0.513 10.424 0.510 0.522
MSE 0.003 0.003 0.003 0.003 0.003 0.003 102.712 0.003 0.0040
$\mathcal{LN}(0.193,1)$ $\Pr(M=M_{2})$ 0.896 0.829 0.952 0.882 0.956 0.954 0.810 0.787 0.875
$\hat{\theta}$ 0.188 0.196 0.187 0.189 0.187 0.191 10.747 0.204 0.187
MSE 0.009 0.008 0.008 0.008 0.009 0.010 108.699 0.013 0.140
$\mathcal{G}\text{a}(2,1)$ $\Pr(M=M_{3})$ 0.952 0.966 0.971 0.984 0.987 0.987 0.175 0.955 0.819
$\hat{\theta}$ 1.000 0.995 0.998 1.004 1.003 1.002 10.789 1.017 1.000
MSE 0.004 0.004 0.004 0.004 0.004 0.004 108.416 0.008 0.275
@@ 4.2.1 Computational Performance Comparison
For $n=100$, the computation times for distance calculations between one observed dataset and one simulated dataset are approximately 0.0018 s for ABC-CvM, 0.0001 s for ABC-Wass, and 0.2685 s for ABC-MMD. For $n=1000$, the times are approximately 0.0040 s for ABC-CvM, 0.0013 s for ABC-Wass, and 0.2738 s for ABC-MMD. Hence, ABC-MMD is the most computationally demanding, with running times more than 2,000 times slower than ABC-Wass and almost 150 times slower than ABC-CvM. This is because standard MMD computation requires pairwise kernel evaluations between all points in both datasets, resulting in $\mathcal{O}(n^{2})$ complexity. In contrast, metrics like the Wasserstein distance can often be computed more efficiently, especially in one dimension. Moreover, MMD relies on kernel functions mapping data into a reproducing kernel Hilbert space (RKHS), adding further computational burden. For comparison, ABC-QDA (Gutmann et al., 2018) takes 0.0529 s to perform QDA on the observed and simulated datasets, making it more computationally demanding than ABC-CvM and ABC-Wass.
@@ 4.3 $g$-and-$k$ Distribution Model Selection
Tables 3 and 4, and Appendix F in the Supplementary Material present results for this example, for sample sizes $n=100$ and $n=1000$. For $n=100$ and when $M_{1}$ is the true model, full data ABC methods yield an average estimated probability around 0.80, which increases to 0.90 when the sample size grows to $n=1000$. ABC-MMD performs slightly worse than the other methods, particularly for the larger sample size, where it still show high variability (see boxplots in the Supplementary Material).
@@
By contrast, NN consistently underperforms. Even for large sample sizes, it tends to assign higher posterior probability to $M_{1}$, leading to frequent misclassification of data generated under $M_{2}$. This indicates that standard neural network architectures may not be well-suited for subtle model discrimination tasks without careful tuning or domain-specific design.
@@
NN struggles to estimate the $g$ parameter, producing highly variable estimates that misleadingly suggest negative skewness. ABC-SA yields estimates comparable to other ABC methods under $M_{1}$, but systematically underestimates $g$ under $M_{2}$. In contrast, ABC-QDA tends to produce similar estimates for $g$ across models $M_{1}$ and $M_{2}$, reducing its ability to distinguish between them.
@@ 6 Conclusion
This Appendix presents additional results for Section 4.1 of the main paper, focusing on the case of known variance. Figure 4 shows the confusion matrices for each problem, while Figures 510 display boxplots of the estimated posterior probabilities of selecting $H_{0}$ across repetitions. Full data ABC methods and ABC-Stat exhibit similar behavior, with high classification accuracy and estimated probabilities of selecting $H_{0}$ approaching zero as the true mean $\theta_{0}$ moves further from zero. When $\theta_{0}$ is small but non-zero, all methods show substantial uncertainty. In comparison, NN struggles across all cases, ABC-QDA consistently selects $H_{0}$ with low uncertainty, and ABC-SA performs better than ABC-QDA and the NN but still shows notable variability in estimated model probabilities.
@@
Figure 11 shows the confusion matrices for the exponential family example, confirming the results reported in the main paper: full data ABC approaches achieve near-perfect classification accuracy, closely matching ABC-Stat. In contrast, ABC-SA and ABC-QDA show lower classification performance, while NN exhibits clear signs of model confusion. The boxplots of the estimated posterior probabilities (Figures 1214) reveal that all methods display a similar variability. Across methods, the gamma model ($M_{3}$) is the easiest to identify (not for NN, which prefers the exponential model to the gamma model even when the data are generated from $M_{3}$), with estimated posterior probabilities close to one in most cases. Models $M_{1}$ and $M_{2}$ are more challenging to recognize, exhibiting higher variability and less consistent model selection.
@@
The confusion matrices (Figure 16 and 17) highlight that classification is considerably more challenging for samples of size $n=100$ than for $n=1000$. Among the methods, ABC-Wass consistently performs best, achieving perfect classification for data generated under $M_{2}$ when $n=1000$. By contrast, ABC-QDA appears to struggle with identifying the correct model even with large samples, while NN tends to favor $M_{1}$, leading to systematic misclassification. Within the full data ABC approaches, ABC-MMD shows comparatively weaker performance.
@@
The boxplots of estimated posterior probabilities of the correct model (Figures 18-21) reveal distinct behaviors across methods. All full data ABC approaches and ABC-Stat exhibit similar performance, with comparable variability and a clear tendency for the estimated probabilities to approach 1 as the sample size increases. ABC-MMD shows the largest variability, particularly when $n=1000$. NN produces extreme probabilities, often close to either 0 or 1, regardless of the true model. ABC-SA exhibits diminished performance for $n=1000$, likely due to the reduced number of summary statistics used as input to limit memory requirements. Finally, ABC-QDA displays the highest variability among all methods, indicating less stable classification.

### v4

@@ 1 Introduction
While deep learning has been applied to related problems such as kinetic model classification (Burés and Larrosa, 2023), such tasks differ conceptually from likelihood-free Bayesian model selection, which explicitly compares the generative models via simulation under each candidate model. Although deep learning classifiers trained with categorical cross-entropy can, in principle, approximate posterior model probabilities, they rely on discriminative approximations that may not remain well-calibrated in overlapping model spaces. ABC, by simulating under each model and comparing to observed data, offers a principled framework for likelihood-free model comparison. In this paper, we evaluate the performance of full data ABC methods for model selection, comparing their effectiveness to summary-based ABC and deep learning classifiers. Our results demonstrate that full data ABC offers a flexible and robust approach to model choice when likelihoods are unavailable.
@@ 4 Simulation Study
The neural network trains on $10^{6}$ datasets (80% for training, 20% for testing) generated using the true parameters, assuming the true model is included in the simulation design. In contrast, ABC simulates from prior distributions and selects samples closest to the observed data using distance metrics, making it more flexible and less assumption-dependent, especially when the true model or parameters are unknown. Training the neural network on datasets generated from the exact model specifications (rather than from a parameter prior) represents an upper-bound or ideal scenario for its performance. In practice, when parameters are drawn from the prior, each model becomes a mixture over many parameter values, increasing within-model variability and making the classification problem more challenging. We conducted a comparison between neural networks trained using fixed parameters and priors on parameters (results not shown here) and found that, apart from the example in Section 4.1, NN exhibited reduced performance across all other examples when trained using parameter priors.
@@ 4.1 Normal Mean Hypothesis Test
Synthetic datasets are generated from the true model $\mathcal{N}(\theta_{0},1)$ with $\theta_{0}\in\{0.0,0.1,0.2,\\ 0.3,0.4,0.5\}$, allowing evaluation of method performance as the true mean departs from $\tilde{\theta}=0$. For each value of $\theta_{0}$, 100 datasets are simulated and analyzed using the seven methods described above for model selection and parameter inference. Under $H_{1}$, the prior is $\theta\sim N(\tilde{\theta},100)$, providing a weakly informative prior; under $H_{0}$, $\theta$ is fixed at $\tilde{\theta}=0$.
@@
We first assess each method’s ability to correctly identify the data-generating model. Performance is summarized using the estimated probability of selecting $H_{0}$ across different values of $\theta_{0}$ (Table 1), and confusion matrices and boxplots of posterior probabilities (Appendix D of the Supplementary Material). ABC methods based on distance metrics correctly identify $H_{0}$ in all cases, indicating good control of false positives. The transition is gradual: for small deviations from zero (e.g., $\theta_{0}=0.1$), the classifier shows limited power, but confidence in selecting $H_{1}$ grows as the true mean increases. When $\theta_{0}\geq 0.4$, $H_{1}$ is selected most of the time, reaching around 95% selection accuracy at $\theta_{0}=0.5$. ABC-SA follows a similar pattern but tends to select $H_{0}$ more often than the full data ABC approaches when the true mean increases. ABC-QDA selects $H_{0}$ in nearly all cases, while NN correctly identifies the model already for smaller values of $\theta_{0}$ (starting from $\theta_{0}=0.2$), although with high variability in posterior probabilities.
We further evaluate posterior inference quality for $\theta$ by computing the MSE between the posterior mean and the true value $\theta_{0}$ across repetitions (Table 1). For each dataset, the posterior mean is computed from the model selected as most probable by each method. Results show that MSE is generally low, with the smallest values observed when $\theta_{0}=0$ and $\theta_{0}=0.5$, consistent with improved classification accuracy for these values. When the true mean is small but non-zero, posterior distributions for all methods retain substantial mass near zero, reflecting limited information in the data to distinguish between models. ABC-SA and ABC-QDA exhibit larger MSE than the distance-based ABC methods, while NN achieves good parameter estimation performance (lowest MSE) for all cases, though this likely reflects an idealized setting where the network is trained on data simulated from the correct models and without uncertainty in $\theta$ under the alternative hypothesis.
@@
These results demonstrate that ABC with distance metrics provides reliable performance for both model selection and parameter estimation, comparable to ABC-Stat using sufficient summary statistics. The comparison with other ABC algorithms in this simple example highlights the limitations of those methods and illustrates the advantages of full data ABC approaches for model selection problems. NN performs well in this simple scenario but becomes unstable in the more complex examples presented below, confirming that the ABC approaches are more robust in realistic inference settings.
@@
$\theta_{0}=0.0$ $\Pr(\theta=0)$ 0.975 0.976 0.975 0.973 0.599 0.976 0.981
$\hat{\theta}$ 0.000 0.000 0.000 0.002 0.042 0.000 0.000
@@
$\theta_{0}=0.1$ $\Pr(\theta=0)$ 0.931 0.933 0.929 0.412 0.925 0.948 0.974
$\hat{\theta}$ 0.007 0.006 0.007 0.010 0.052 0.000 0.000
@@
$\theta_{0}=0.2$ $\Pr(\theta=0)$ 0.788 0.796 0.789 0.760 0.261 0.847 0.955
$\hat{\theta}$ 0.046 0.042 0.042 0.048 0.155 0.010 0.000
MSE 0.037 0.037 0.038 0.038 0.013 0.037 0.040
$\theta_{0}=0.3$ $\Pr(\theta=0)$ 0.522 0.543 0.519 0.459 0.084 0.671 0.934
$\hat{\theta}$ 0.147 0.144 0.149 0.168 0.273 0.141 0.000
MSE 0.062 0.061 0.060 0.056 0.004 0.096 0.090
$\theta_{0}=0.4$ $\Pr(\theta=0)$ 0.210 0.236 0.203 0.157 0.005 0.450 0.902
$\hat{\theta}$ 0.344 0.335 0.345 0.364 0.404 0.285 0.000
MSE 0.039 0.041 0.038 0.034 0.000 0.125 0.160
$\theta_{0}=0.5$ $\Pr(\theta=0)$ 0.056 0.067 0.050 0.031 0.014 0.263 0.870
$\hat{\theta}$ 0.494 0.489 0.497 0.506 0.484 0.359 0.000
MSE 0.022 0.025 0.020 0.016 0.003 0.125 0.250
@@ 4.2 Exponential Family Model Selection
NN is the method that shows the lowest accuracy of model selection: it correctly identifies the exponential model in only 55% of cases, the lognormal in 65%, and the gamma in around 97%. The lognormal model was often misclassified as exponential, suggesting a structural limitation in distinguishing these distributions.
Parameter estimation results (Table 2) mirror these findings. For distance-based ABC methods, MSEs are consistently low when the correct model is selected, comparable to ABC-Stat. NN produces slightly more biased estimates than the ABC methods, in particular for datasets generated from the exponential and the lognormal models, reflecting its sensitivity to misclassification errors. ABC-SA yields slightly higher estimation errors. Finally, ABC-QDA has the largest MSE values across all experiments and ABC methods.
@@
$\mathcal{E}\text{xp}(0.5)$ $\Pr(M=M_{1})$ 0.883 0.836 0.953 0.850 0.948 0.952 0.554 0.772 0.922
$\hat{\theta}$ 0.512 0.506 0.511 0.515 0.514 0.513 0.454 0.510 0.522
MSE 0.003 0.003 0.003 0.003 0.003 0.003 0.021 0.003 0.0040
$\mathcal{LN}(0.193,1)$ $\Pr(M=M_{2})$ 0.896 0.829 0.952 0.882 0.956 0.954 0.653 0.787 0.875
$\hat{\theta}$ 0.188 0.196 0.187 0.189 0.187 0.191 0.337 0.204 0.187
MSE 0.009 0.008 0.008 0.008 0.009 0.010 0.034 0.013 0.140
$\mathcal{G}\text{a}(2,1)$ $\Pr(M=M_{3})$ 0.952 0.966 0.971 0.984 0.987 0.987 0.967 0.955 0.819
$\hat{\theta}$ 1.000 0.995 0.998 1.004 1.003 1.002 0.957 1.017 1.000
MSE 0.004 0.004 0.004 0.004 0.004 0.004 0.008 0.008 0.275
@@ 4.2.1 Computational Performance Comparison
For $n=100$, the computation times for distance calculations between one observed dataset and one simulated dataset are approximately 0.0018 s for ABC-CvM, 0.0001 s for ABC-Wass, and 0.2685 s for ABC-MMD. For $n=1000$, the times are approximately 0.0040 s for ABC-CvM, 0.0013 s for ABC-Wass, and 0.2738 s for ABC-MMD. As a comparison, ABC-Stat takes about 0.0053 s for $n=100$ and 0.0582 s for $n=1000$, showing similar computational cost to ABC-CvM for the examples considered in this paper. Hence, ABC-MMD is the most computationally demanding, with running times more than 2,000 times slower than ABC-Wass and almost 150 times slower than ABC-CvM. This is because standard MMD computation requires pairwise kernel evaluations between all points in both datasets, resulting in $\mathcal{O}(n^{2})$ complexity. In contrast, metrics like the Wasserstein distance can often be computed more efficiently, especially in one dimension. Moreover, MMD relies on kernel functions mapping data into a reproducing kernel Hilbert space (RKHS), adding further computational burden. For comparison, ABC-QDA (Gutmann et al., 2018) takes 0.0529 s to perform QDA on the observed and simulated datasets, making it more computationally demanding than ABC-CvM and ABC-Wass.
The computational efficiency of ABC-Stat depends heavily on the number and nature of summary statistics used to represent the data: simple, low-dimensional summaries make it comparable to ABC-CvM, while more complex or high-dimensional statistics can lead to a substantial increase in runtime. However, for the examples presented in this paper, where the chosen summaries are simple and low-dimensional, the computation performance of ABC-Stat is very fast.
For completeness, we also measured the computational cost of ABC-SA, which involves an initial summary-statistic selection step followed by ABC rejection. ABC-SA is not directly comparable with the other methods, because it was run on a different software and directly on the full set of input summary statistics. When implemented using the selectsumm() function from the abctools package in R with $10^{6}$ simulated datasets and all observations as input summary statistics, each full call requires 540.7 s for $n=100$ and 1620.9 s for $n=1000$. It should be noted, however, that this comparison is not directly equivalent to the other methods (which are Python-based), since the R implementation performs the entire selection and rejection step jointly across all simulations, and involves additional overhead from the interpreted environment. The runtime of ABC-SA scales approximately linearly with the number of simulations and with the dimension of the summary vector, since the selection step must evaluate all simulations jointly. Consequently, ABC-SA remains efficient for moderate-sized problems but may become slower for very high-dimensional summaries or larger simulation budgets.
@@ 4.3 $g$-and-$k$ Distribution Model Selection
Tables 3 and 4, and Appendix F in the Supplementary Material present results for this example, for sample sizes $n=100$ and $n=1000$. For $n=100$ and when $M_{1}$ is the true model, full data ABC methods yield an average estimated probability around 0.80, which increases to 0.90 when the sample size grows to $n=1000$. ABC-MMD performs slightly worse than the other methods, particularly for the larger sample size, where it still shows high variability (see boxplots in the Supplementary Material).
@@
NN achieves the best identification performance for model $M_{2}$, but it struggles to correctly recognize data generated under $M_{1}$. The estimation of the $g$ parameter shows the lowest MSE when $M_{2}$ is true but the highest MSE under $M_{1}$. This pattern remains consistent across all sample sizes, indicating that increased data availability does not mitigate this bias. NN therefore performs well when the underlying asymmetry is identifiable, but it remains sensitive to overlapping model features. This highlights that while neural network classifiers can capture complex dependencies, their performance depends on how well models are separated.
@@ 6 Conclusion
This Appendix presents additional results for Section 4.1 of the main paper, focusing on the case of known variance. Figure 4 shows the confusion matrices for each problem, while Figures 510 display boxplots of the estimated posterior probabilities of selecting $H_{0}$ across repetitions. Full data ABC methods and ABC-Stat exhibit similar behavior, with high classification accuracy and estimated probabilities of selecting $H_{0}$ approaching zero as the true mean $\theta_{0}$ moves further from zero. When $\theta_{0}$ is small but non-zero, all methods show substantial uncertainty. In comparison, ABC-QDA consistently selects $H_{0}$ with low uncertainty, and ABC-SA performs better than ABC-QDA and the NN but still shows notable variability in estimated model probabilities. NN performs well across all settings, displaying the highest uncertainty in the most challenging case ($\theta_{0}=0.1$), but then clearly decreasing the estimated probability of selecting $H_{0}$ as $\theta_{0}$ increases.
@@
Figure 11 shows the confusion matrices for the exponential family example, confirming the results reported in the main paper: full data ABC approaches achieve near-perfect classification accuracy, closely matching ABC-Stat. In contrast, ABC-SA and ABC-QDA show lower classification performance, while NN displays confusion between the exponential and lognormal models and exhibits the highest variability in the estimated probabilities. The boxplots of the estimated posterior probabilities (Figures 1214) reveal that all methods display a similar variability. Across methods, the gamma model ($M_{3}$) is the easiest to identify, with estimated posterior probabilities close to one in most cases. Models $M_{1}$ and $M_{2}$ are more challenging to recognize, exhibiting higher variability and less consistent model selection.
@@
The confusion matrices (Figure 16 and 17) highlight that classification is considerably more challenging for samples of size $n=100$ than for $n=1000$. Among the methods, ABC-Wass consistently performs best, achieving perfect classification for data generated under $M_{2}$ when $n=1000$. By contrast, ABC-QDA appears to struggle with identifying the correct model even with large samples, while NN tends to favor $M_{1}$, leading to systematic misclassification when data are generated from $M_{0}$. Within the full data ABC approaches, ABC-MMD shows comparatively weaker performance.
@@
The boxplots of estimated posterior probabilities of the correct model (Figures 18-21) reveal distinct behaviors across methods. All full data ABC approaches and ABC-Stat exhibit similar performance, with comparable variability and a clear tendency for the estimated probabilities to approach 1 as the sample size increases. ABC-MMD shows the largest variability, particularly when $n=1000$. NN produces posterior probabilities that are typically very close to one when the data are generated from $M_{2}$, but exhibit substantial variability when the data are generated from $M_{1}$, reflecting uncertainty and instability in model discrimination. ABC-SA exhibits diminished performance for $n=1000$, likely due to the reduced number of summary statistics used as input to limit memory requirements. Finally, ABC-QDA displays the highest variability among all methods, indicating less stable classification.